OpenStudy (anonymous):

Write the equation of the line that is parallel to the line y = 2x + 2 and passes through the point (5, 3). A. y = 2x − 7 B. y = −1/2x − 2 C. y = 2x − 2 D. y = −1/2x − 7

OpenStudy (anonymous):

Use:\[y - y _{1} = m(x - x _{1})\]where "m" is your slope value from your original equation that is in the form y = mx + b

OpenStudy (anonymous):

(5, 3) is your point (x1, y1). So, now you can substitute and simplify.

OpenStudy (anonymous):

wait, what?..

OpenStudy (anonymous):

what?